OpenStudy (anonymous):

simplify ln(x^2+5x) =2ln(x+1)

OpenStudy (anonymous):

you mean solve for x?

OpenStudy (anonymous):

right sorry my bad

OpenStudy (anonymous):

start with \[\ln(x^2+5x)=\ln((x+1)^2)\] and then since log is a one to one function this means \[x^2+5x=(x+1)^2\] solve the quadratic equation

OpenStudy (anonymous):

thank you, can you help me also with thise question below: thanks! http://openstudy.com/study?F5769872436922CDHW5=_#/updates/4f19c38be4b04992dd225045